OpenStudy (gabylovesyou):

Which of the following is equivalent to log16 8? 2 1.248 4 over 3 3 over 4

OpenStudy (gabylovesyou):

\[\log _{16}8\]

satellite73 (satellite73):

got a calcualtor?

OpenStudy (gabylovesyou):

yes

satellite73 (satellite73):

type in this \[\log(8)\div \log(16)\]

satellite73 (satellite73):

this is a non calculator way to do it as well, if you can figure out how to solve \[16^x=8\]

OpenStudy (gabylovesyou):

3/4

satellite73 (satellite73):

yup`